OpenStudy (anonymous):

how do I resolve into partial fractions: 3x^2-8x+1/(x-4)(x^2+1)

OpenStudy (candi5a18):

synthetic division first

hartnn (hartnn):

I think synthetic division is not required here, the degree of numerator is already less than the denominator. so we do this: \(\Large \dfrac{3x^2-8x+1}{(x-4)(x^2+1)}=\dfrac{A}{x-4}+\dfrac{Bx+C}{x^2+1}\) let me know if you need help in finding A,B,C. @salii99

hartnn (hartnn):

hint: make common denominator on right side, and then compare co-efficients of the numerators on both sides.